# Pixel Is Not A Barrier: An Effective Evasion Attack for Pixel-Domain Diffusion Models

### v1

@@ 4.4 Alternating Optimization for Adversarial Image
We solve the constrained optimization problems via alternating optimization to craft the adversarial images, detailed optimization loop of AtkPDM+ is provided in Algorithm  1. AtkPDM algorithm and the derivation of the alternating optimization are provided in Appendix.
@@ 4.5 Latent Optimization via Pretrained-VAE
Previous works suggest that diffusion models have a strong capability of resisting adversarial perturbations (Xue and Chen 2024), making them hard to attack via pixel-domain optimization. Moreover, they are even considered good purifiers of adversarial perturbations (Nie et al. 2022). Here we propose a latent optimization strategy that crafts the “perturbation” in latent space. We adopt a pre-trained Variational Autoencoder (VAE)  (Kingma and Welling 2014) to convert images to their latent space, and the gradients will be used to update the latent, after N iterations or losses converge, we decode back via decoder $\mathcal{D}$ to pixel domain as our final protected image. The motivation for adopting VAE is inspired by MPGD (He et al. 2024). This strategy is effective for crafting a robust adversarial image against pixel-domain diffusion models while also better preserving the protection quality rather than only incorporating fidelity constraints. The constraint optimization thereby becomes:
$\max_{\mathbf{z}^{\operatornamewithlimits{adv}}}\mathbb{E}_{t,\mathbf{x}_{t}|\mathbf{x},\mathbf{x}_{t}^{\operatornamewithlimits{adv}}|\mathcal{D}(\mathbf{z}^{\operatornamewithlimits{adv}})}\left[-\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}_{t}^{\operatornamewithlimits{adv}})\right]$ (7)
subject to $\text{subject to }\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathcal{D}(\mathbf{z}^{\operatornamewithlimits{adv}}))\leq\delta.$
Detailed latent optimization loop is provided in Algorithm 1.
@@
1:  Input: Image to be protected $\mathbf{x}$, attack budget $\delta>0$, step size $\gamma_{1},\gamma_{2}>0$, VAE encoder $\mathcal{E}$, and VAE decoder $\mathcal{D}$
@@
7:     Sample noise: $\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
8:     Compute original noisy sample: $\mathbf{x}_{t}\leftarrow\mathcal{F}(\mathbf{x},t,\epsilon_{1})$
9:     Compute adversarial noisy sample: $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}\leftarrow\mathcal{F}(\mathbf{x}^{\operatornamewithlimits{adv}},t,\epsilon_{2})$
10:     Update $\mathbf{z}^{\operatornamewithlimits{adv}}$ by Gradient Descent: $\mathbf{z}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{z}^{\operatornamewithlimits{adv}}-\gamma_{1}\operatornamewithlimits{sign}(\nabla_{\mathbf{z}^{\operatornamewithlimits{adv}}}\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}^{\operatornamewithlimits{adv}}_{t}))$
@@
12:         $\mathbf{z}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{z}^{\operatornamewithlimits{adv}}-\gamma_{2}\nabla_{\mathbf{z}^{\operatornamewithlimits{adv}}}\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathcal{D}(\mathbf{z}^{\operatornamewithlimits{adv}}))$
@@ 5 Experiment Results
In this section, we examine the attack effectiveness and robustness of our approach under extensive settings.
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA-Score $\downarrow$
Church PGAscent 0.37 $\pm$ 0.09 28.17 $\pm$ 0.22 0.73 $\pm$ 0.16 0.89 $\pm$ 0.05 31.06 $\pm$ 1.94 0.17 $\pm$ 0.09 0.93 $\pm$ 0.04
@@
Table 1: Quantitative Results in attacking different PDMs. The best is marked in red and the second best is marked in blue. Errors denote one standard deviation of all images in our test datasets.
@@
SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA-Score $\downarrow$
@@
Table 2: Quantitative results of our adversarial images against defense methods. Both Crop-and-Resize and JPEG Compression fail to defend our attack. “None” indicates no defense is applied, as the baseline for comparison.
@@ 5.1 Experiment Settings
We conduct all our experiments in white box settings and examine the effectiveness of our attacks using SDEdit (Meng et al. 2021). For the Variational Autoencoder (VAE)  (Kingma and Welling 2014) in our AtkPDM+, we utilize the VAE provided by StableDiffusion V1.5  (Rombach et al. 2022). We run all of our experiments with 300 optimization steps, which is empirically determined, to balance attacking effectiveness and image protection quality with reasonable speed. Other loss parameters and running time are provided in the Appendix. The implementation is built on the Diffusers library. All the experiments are conducted with a single Nvidia Tesla V100 GPU.
@@
We test our approach on PDMs with three open-source checkpoints on HuggingFace, specifically “google/ddpm-ema-church-256”, ”google/ddpm-cat-256” and “google/ddpm-ema-celebahq-256”. For the results reported in Table 1, we run 30 images for each victim model. Additionally, for generalizability in practical scenarios, we synthesize the data with half randomly from the originally trained dataset and another half from randomly crawled with keywords from the Internet.
@@
To the best of our knowledge, previous methods have mainly focused on LDMs, and effective PDM attacks have not yet been developed, however, we still implement Projected Gradient Ascent (PGAscent) with their proposed semantic loss by (Salman et al. 2023; Liang et al. 2023; Liang and Wu 2023; Xue et al. 2024b). Notably, Diff-Protect (Xue et al. 2024b) proposed to minimize the semantic loss is surprisingly better than maximizing the semantic loss, we also adopted this method in attacking PDMs and denote as Diff-Protect. To quantify the adversarial image visual quality, we adopted Structural Similarity (SSIM)  (Wang et al. 2004), Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity (LPIPS)  (Zhang et al. 2018). We also inherit these three metrics, but negatively to quantify the effectiveness of our attack. We also adopted Image Alignment Score (IA-Score) (Kumari et al. 2023) that leverages CLIP (Radford et al. 2021) to calculate the cosine similarity of image encoder features. In distinguishing from previous methods, to more faithfully reflect the attack effectiveness, we fix the same seed of the random generator when generating clean and adversarial samples, then calculate the scores based on the paired samples.
5.2 Attack Effectiveness on PDMs
As quantitatively reported in Table 1 and qualitative results in Figure 4, compared to previous PGD-based methods incorporating semantic loss, i.e., negative training loss of diffusion models, our method exhibits superior performance in both adversarial image quality and attacking effectiveness. And our reported figures has generally stable as reflected in lower standard deviation. It is worth noting that even if the adversarial image qualities of the PGD-based methods are far worse than ours, their attacking effectiveness still falls short, suggesting that PDMs are robust against traditional perturbation methods, this finding is also aligned with previous works (Xue et al. 2024b; Xue and Chen 2024). For AtkPDM+, combined with our latent optimization strategy, the adversarial image quality has enhanced while slightly affecting the attacking effectiveness, still outperforming the previous methods.
@@ 5.2 Attack Effectiveness on PDMs
SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA-Score $\downarrow$
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA-Score $\downarrow$
@@
$\mathcal{L}_{\text{attack}}$ + $\mathcal{L}_{\text{fidelity}}$ (Ours) 0.75 $\pm$ 0.03 28.22 $\pm$ 0.10 0.26 $\pm$ 0.04 0.75 $\pm$ 0.04 29.61 $\pm$ 0.23 0.40 $\pm$ 0.05 0.76 $\pm$ 0.06
$\mathcal{L}_{\text{attack}}$ + $\mathcal{L}_{\text{fidelity}}$ (Ours) 0.81 $\pm$ 0.03 28.64 $\pm$ 0.19 0.13 $\pm$ 0.02 0.79 $\pm$ 0.04 30.05 $\pm$ 0.47 0.33 $\pm$ 0.07 0.81 $\pm$ 0.06
Table 4: Quantitative results of ablation study. The best is in bold and the second best is underlined. Errors denote one standard deviation of all images in our test datasets.
5.3 Against Defense Methods
We examine the robustness of our approach against two widely recognized and effective defense methods for defending against adversarial attacks as reported in Table 2.
@@ 5.3 Against Defense Methods
Noted by Diff-Protect, crop and resize is simple yet the most effective defense method against their attacks on LDMs. We also test our method against this defense using their settings, i.e., cropping 20% of the adversarial image and then resizing it to its original dimensions.
@@
Sandoval-Segura et al. (Sandoval-Segura, Geiping, and Goldstein 2023) demonstrated that JPEG compression is a simple yet effective adversarial defense method. In our experiments, we implement the JPEG compression at a quality setting of 25%. The quantitative results in Table 2 demonstrate that our method is robust against these two defense methods, with four of the metrics listed in Table 2 are not worse than no defenses. Surprisingly, these defense methods even make the adversarial image more effective than cases without defense.
5.4 Black Box Transferability
We craft adversarial images with the proxy model, “google/ddpm-ema-church-256”, in white-box settings and test their transferability to another “google/ddpm-bedroom-256” model for black-box attacks. Under identical validation settings, Table 3 reveals only a slight decrease in attack effectiveness metrics, indicating successful black-box transferability.
@@ 5.5 Effectiveness of Latent Optimization via VAE
We first incorporate our VAE latent optimization strategy in the previous semantic-loss-based PGAscent. From Table 4, without using $\mathcal{L}_{\text{fidelity}}$, latent optimization has significantly enhanced the adversarial image quality and even slightly improved the attacking effectiveness. Adopting latent optimization in our approach enhances visual quality with a negligible decrease in attacking effectiveness. Surprisingly, incorporating our $\mathcal{L}_{\text{fidelity}}$ with current PGD-based method will drastically decrease the adversarial image quality despite its attack performing better than ours. This may be due to different constrained optimization problem settings.
@@ 6 Conclusion
In this paper, we present the first framework designed to effectively protect images manipulation by Pixel-domain Diffusion Models (PDMs). We demonstrate that while denoising UNets appear robust to conventional PGD-based attacks, their feature space remains vulnerable to attack. Our proposed feature attacking loss exploit the vulnerabilities to empower adversarial images to mislead PDMs, thereby producing low-quality output images. We approach this image protection problem as a constrained optimization problem, solving it through alternating optimization. Additionally, our latent optimization strategy via VAE enhances the naturalness of our adversarial images. Extensive experiments validate the efficacy of our method, achieving state-of-the-art performance in attacking PDMs.
Supplementary Material

### v3

@@ 3.4 Alternating Optimization for Adversarial Image
We solve the constrained optimization problem via alternating optimization to craft the adversarial images, detailed optimization loop of AtkPDM+ is provided in Algorithm  1. To maximize the $\mathcal{L}_{\text{attack}}$, we take the negative $\mathcal{L}_{\text{attack}}$ and perform gradient descent. AtkPDM algorithm and the derivation of the alternating optimization are provided in Appendix.
@@ 3.5 Latent Optimization via Pretrained-VAE
Previous works suggest that diffusion models have a strong capability against adversarial perturbations (Xue and Chen 2024), making them hard to be attacked via pixel-domain optimization. Moreover, they are even considered as good purifiers of adversarial perturbations (Nie et al. 2022).
Here, we propose a strategy that crafts the perturbation in the latent space of the pre-trained Variational Autoencoder (VAE)  (Kingma and Welling 2014), and the gradients are used to update the latent. After $N$ iterations or losses converge, we decode back via the decoder $\mathcal{D}$ to pixel domain as our final adversarial image. The motivation for adopting VAE is inspired by MPGD (He et al. 2024). This strategy is effective for crafting a robust adversarial image against pixel-domain diffusion models while also better preserving the adversarial image quality rather than only incorporating fidelity constraints. Note that, ideally, manifold preservation is guaranteed when using perfect VAE. In practice, we use the best available LDM’s VAE agnostic to the victim model as our latent optimization VAE. Detailed latent optimization loop is provided in Algorithm 1.
@@
1:  Input: Image to be protected $\mathbf{x}$, attack budget $\delta>0$, step size $\gamma_{\text{attack}},\gamma_{\text{fidelity}}>0$, VAE encoder $\mathcal{E}$, and VAE decoder $\mathcal{D}$
@@
7:     Sample noise: $\epsilon,\epsilon^{\operatornamewithlimits{adv}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
8:     Compute original noisy sample: $\mathbf{x}_{t}\leftarrow\mathcal{F}(\mathbf{x},t,\epsilon)$
9:     Compute adversarial noisy sample: $\mathbf{x}^{\operatornamewithlimits{adv}}_{t}\leftarrow\mathcal{F}(\mathbf{x}^{\operatornamewithlimits{adv}},t,\epsilon^{\operatornamewithlimits{adv}})$
10:     Update $\mathbf{z}^{\operatornamewithlimits{adv}}$ by Gradient Descent: $\mathbf{z}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{z}^{\operatornamewithlimits{adv}}-\gamma_{\text{attack}}\operatornamewithlimits{sign}(\nabla_{\mathbf{z}^{\operatornamewithlimits{adv}}}(-\mathcal{L}_{\text{attack}}(\mathbf{x}_{t},\mathbf{x}^{\operatornamewithlimits{adv}}_{t})))$
@@
12:         $\mathbf{z}^{\operatornamewithlimits{adv}}\leftarrow\mathbf{z}^{\operatornamewithlimits{adv}}-\gamma_{\text{fidelity}}\nabla_{\mathbf{z}^{\operatornamewithlimits{adv}}}\mathcal{L}_{\text{fidelity}}(\mathbf{x},\mathcal{D}(\mathbf{z}^{\operatornamewithlimits{adv}}))$
@@ 4 Experiment Results
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA $\downarrow$
Church AdvDM 0.37 $\pm$ 0.09 28.17 $\pm$ 0.22 0.73 $\pm$ 0.16 0.89 $\pm$ 0.05 31.06 $\pm$ 1.94 0.17 $\pm$ 0.09 0.93 $\pm$ 0.04
@@
Table 1: Quantitative results in attacking different unconditional PDMs. The best is marked in bold and the second best is underlined. Errors denote one standard deviation of all images in our test datasets.
Methods Adversarial Image Quality Attacking Effectiveness
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA $\downarrow$
Diff-Protect 0.47 $\pm$ 0.08 27.96 $\pm$ 0.08 0.46 $\pm$ 0.05 0.49 $\pm$ 0.10 28.13 $\pm$ 0.15 0.36 $\pm$ 0.10 0.79 $\pm$ 0.06
AtkPDM+ (Ours) 0.79 $\pm$ 0.06 28.48 $\pm$ 0.33 0.06 $\pm$ 0.02 0.72 $\pm$ 0.10 28.50 $\pm$ 0.48 0.10 $\pm$ 0.04 0.86 $\pm$ 0.08
Table 2: Quantitative results in attacking conditional PDM DeepFloyd IF. The best is marked in bold and the second best is underlined. Errors denote one standard deviation of all images in our test datasets.
@@
SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA $\downarrow$
LDM-Pure 0.78 29.84 0.35 0.80
@@
Table 3: Quantitative results of our adversarial images against defense methods. LDM-Pure, Crop-and-Resize, and JPEG Compression fail to defend our attack. “None” indicates no defense is applied, as the baseline for comparison.
@@ 4.1 Experiment Settings
We conduct all our experiments in white box settings and examine the effectiveness of our attacks using SDEdit (Meng et al. 2021). For the VAE (Kingma and Welling 2014) in our AtkPDM+, we utilize the one provided by StableDiffusion v1.5 (Rombach et al. 2022). We run all of our experiments with 300 optimization steps, which empirically determined, balancing attacking effectiveness and adversarial image quality with a reasonable speed. Other loss parameters and running time are provided in the Appendix. The implementation is built on the Diffusers library (von Platen et al. 2022). All the experiments are conducted with a single Nvidia Tesla V100 GPU.
@@
We test our approach on PDMs with three open-source checkpoints on HuggingFace, specifically “google/ddpm-ema-church-256”, “google/ddpm-cat-256” and “google/ddpm-ema-celebahq-256”. For the results reported in Table 1, we run 30 images for each victim model. Additionally, for generalizability in practical scenarios, we synthesize the data with half randomly selected from the originally trained dataset and another half from randomly crawled with keywords from the Internet.
@@
To the best of our knowledge, the previous methods have mainly focused on LDMs, and effective PDM attacks have not yet been developed, however, we still implement AdvDM  (Liang et al. 2023) with the proposed semantic loss by (Salman et al. 2023; Liang et al. 2023; Liang and Wu 2023; Xue et al. 2023) for comparison. Notably, Diff-Protect (Xue et al. 2023) proposed to minimize the semantic loss and is counterintuitively better than maximizing the semantic loss. We also adopt this method in attacking PDMs. To quantify the adversarial image visual quality, we adopt Structural Similarity (SSIM) (Wang et al. 2004), Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity (LPIPS)  (Zhang et al. 2018) as the evaluation metrics but negatively quantify the effectiveness of our attack. We also adopt the Image Alignment Score (IA) (Kumari et al. 2023) that leverages CLIP (Radford et al. 2021) to calculate the cosine similarity between two image encoder features. In distinguishing from the previous methods, to more faithfully reflect the attacking effectiveness, we fix the same seed of the random generator when generating clean and adversarial samples, then calculating the scores based on the paired samples.
4.2 Attacking Effectiveness on PDMs
As quantitatively reported in Table 1 and qualitative results in Figure 4, compared to the previous PGD-based methods incorporating semantic loss, i.e., negative training loss of diffusion models, our method exhibits superior performance in both adversarial image quality and attacking effectiveness. In addition, our reported numbers are generally stable, as reflected in lower standard deviation. It is worth noting that even if the adversarial image qualities of the PGD-based methods are far worse than ours, their attacking effectiveness still falls short, suggesting that PDMs are robust against traditional perturbation methods. This finding is also aligned with previous works (Xue et al. 2023; Xue and Chen 2024). For AtkPDM+, combined with our latent optimization strategy, the adversarial image quality has been enhanced while slightly affecting the attacking effectiveness, still outperforming the previous methods. Besides unconditional PDMs, we also compare with the previous best method Diff-Protect against a conditional PDM DeepFloyd IF (at StabilityAI 2023), reported in Table  2. Although the attacking effectiveness of Diff-Protect seems better than ours, this may be due to their adversarial image quality being severely corrupted during the attack. Hence, it cannot fulfill our two objectives simultaneously. In addition, our framework is extensible to attack LDMs, please refer to Appendix provided in the project page.
@@ 4.2 Attacking Effectiveness on PDMs
SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA $\downarrow$
@@
Figure 4: Qualitative results compared to the previous methods. Our adversarial images can effectively corrupt the edited results without significant fidelity decrease. The same column shares the same random seed for fair comparisons.
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\downarrow$ PSNR $\downarrow$ LPIPS $\uparrow$ IA $\downarrow$
@@
$\mathcal{L}_{\text{attack}}$ + $\mathcal{L}_{\text{fidelity}}$ 0.75 $\pm$ 0.03 28.22 $\pm$ 0.10 0.26 $\pm$ 0.04 0.75 $\pm$ 0.04 29.61 $\pm$ 0.23 0.40 $\pm$ 0.05 0.76 $\pm$ 0.06
$\mathcal{L}_{\text{attack}}$ + $\mathcal{L}_{\text{fidelity}}$ 0.81 $\pm$ 0.03 28.64 $\pm$ 0.19 0.13 $\pm$ 0.02 0.79 $\pm$ 0.04 30.05 $\pm$ 0.47 0.33 $\pm$ 0.07 0.81 $\pm$ 0.06
Table 5: Quantitative results of ablation study. The best is marked in bold and the second best is underlined. Errors denote one standard deviation of all images in our test datasets.
4.3 Black Box Transferability
We craft adversarial images with the proxy model, “google/ddpm-ema-church-256”, in white-box settings and test their transferability against “google/ddpm-bedroom-256” model as black-box attacks. Under identical validation settings, Table 4 reveals only a slight decrease in attack effectiveness metrics, suggesting black-box transferability.
4.4 Robustness Against Defense Methods
We examine the robustness of our approach against three widely recognized and effective adversarial defense methods. The quantitative results in Table 3 demonstrate that our method is robust against these three defense methods, with four metrics listed in Table 3 not worse than no defenses. Surprisingly, these defense methods even make the adversarial image more effective than cases without defense. We provide the implementation details of each defense method in the following sections.
LDM Purification.
Nie at. al. proposed DiffPure  (Nie et al. 2022) that leverages a pre-trained Diffusion Model to purify adversarial images targeting classifier models to defend effectively. The purification process is essentially an unconditional SDEdit process with small forward $t$. Here, we use a pre-trained LDM (StableDiffusion v1.5) and $t=100$ to purify our adversarial image as a defense method.
@@ 4.4 Robustness Against Defense Methods
Noted by Diff-Protect, “crop and resize” is a simple yet the most effective defense method against their attacks on LDMs. We test our method against this defense using their settings, i.e., cropping 20% of the adversarial image and resizing it to its original dimensions.
@@
Sandoval-Segura et al. (Sandoval-Segura, Geiping, and Goldstein 2023) demonstrated that JPEG compression is a simple yet effective adversarial defense method. In our experiments, we implement the JPEG compression at a quality setting of 25%.
@@ 4.5 Effectiveness of Latent Optimization via VAE
We first incorporate our VAE latent optimization strategy in the previous semantic-loss-based methods. From Table 5, without using $\mathcal{L}_{\text{fidelity}}$, latent optimization has significantly enhanced the adversarial image quality and even slightly improved the attacking effectiveness. Adopting latent optimization in our approach enhances visual quality with a negligible decrease in attacking effectiveness. Surprisingly, incorporating our $\mathcal{L}_{\text{fidelity}}$ with current PGD-based method will drastically decrease the adversarial image quality despite its attack performing better than ours. This may be due to different constrained optimization problem settings.
@@ 5 Conclusion
This paper presents the first framework to protect against image manipulation by Pixel-domain Diffusion Models (PDMs). While denoising UNets withstand traditional PGD attacks, their feature space remains vulnerable. Our feature attacking loss exploits these vulnerabilities, generating adversarial images that mislead PDMs, resulting in corrupted output. We approach this image protection problem as a constrained optimization problem, solving it through alternating optimization. Furthermore, our latent optimization strategy via VAE enhances the naturalness of our adversarial images. Extensive experiments validate the efficacy of our method, achieving state-of-the-art performance in attacking PDMs.
Acknowledgements
This research is supported by National Science and Technology Council, Taiwan (R.O.C) under the grant numbers NSTC-113-2634-F-002-007, NSTC-112-$2$-E-001-001-MY2, NSTC-113-2634-F-001-002-MBK, NSTC-113-2221-E-002-201, and Academia Sinica under the grant number of AS-CDA-110-M09. We thank to National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.
Appendix